# The Too Visible QCD Axion Luca Di Luzioa, Michele Redib, Alessandro Strumiac, Andrea Tesib, Arsenii V. Titovd,a

### v1

@@ 3.4 Model with axion coupled to the strange quark
$m_{\varphi}=\left(\frac{B_{0}f_{\pi}^{2}}{2m_{s}}\right)^{1/2}\kappa_{s}\approx 0.36~\text{GeV}\,\kappa_{s}.$ (32)
@@
Extra tentative models with multiple PQ scalars are discussed in appendix A. One might even consider the possibility of adding a PQ-neutral scalar just to make the $\pi^{0}$ heavier. In section 4 we discuss the general reason that prevents getting a viable model.
@@ 3.5 Loop corrections to pion masses
All results receive corrections that get large if $\kappa$ is large. Here we attempt an estimate. One-loop corrections to pion masses directly affect the dominant diagonal entries of the meson mass matrix in eq. (17). So there could be a controllable regime where the one-loop corrections cancel the too large tree-level correction to $m_{\pi^{0}}-m_{\pi^{\pm}}$, arising from off-diagonal entries in the meson mass matrix. We here heuristically derive the pion mass splitting by computing the renormalization of the meson masses in the chiral Lagrangian effective theory, including $\varphi$. In addition to isospin-symmetric vertices, it contains a $\sigma\pi^{0}\eta^{(\prime)}$ vertex that only affects the $\pi^{0}$ mass. The linear coupling of $\varphi$ to QCD mesons can be written in terms of $\varphi$-dependent meson mass matrix, obtaining the 1-loop contribution as
@@
where latin indices label QCD mesons. The Euclidean integral is positive. The integral is logarithmically UV divergent. More importantly, we cannot apply the chiral Lagrangian at large $Q\gtrsim$ GeV. So we insert in the above expression a cut-off at $\Lambda\sim m_{\rho}\sim{\rm GeV}$. Assuming $m_{\varphi}\gtrsim\Lambda$ and neglecting for simplicity off-diagonal terms in the mass matrix, we find
@@
This splitting, when extrapolated to $m_{\varphi}>\Lambda$ agrees, including the sign, with the result of the tree level matching eq. (24) upon identifying $\Lambda=4\pi f_{\pi}$. This however should not be considered as a new contribution to the splitting but rather as an estimate of the uncertainty.
@@ 4 Heavy axion effective chiral theory
We here focus on the regime where the new PQ scalars ($a$ and $\sigma$ in the minimal model) are heavy enough that they could be hidden among QCD resonances with GeV mass. This regime needs $n_{s}=0$ and $n_{u,d}\leq 1$ in the language of the previous section. We here exploit the (mild) separation of scales
@@
to integrate out the PQ scalars, obtaining a modified chiral Lagrangian for the light mesons. Namely we capture the effect of the heavy field by matching the full theory of eq. (3) at $m_{\varphi}$ to a new chiral Lagrangian augmented by operators — invariant under PQ — generated by the new scalar dynamics. This allows to exploit the symmetries to constrain the form of extra effects, including higher loop effects. The low energy degrees of freedom are again the $\Pi$ fields of eq. (9). The $\eta^{\prime}$ is again kept in the effective field theory (EFT) because of its special role for the PQ symmetry. The meson interactions differ from the usual chiral Lagrangian, and even include different isospin-breaking terms. Indeed, at the matching scale, formal invariants under the full symmetry now involve two spurions:
@@
The first term, $\mathscr{L}_{\rm QCD}(\tilde{M}_{q})$, is the chiral Lagrangian with quark masses given by $\tilde{M}_{q}$. Like in eq. (7), it includes the term ${f_{\pi}^{2}B_{0}}\operatorname{tr}[\tilde{M}_{q}U^{\dagger}+\tilde{M}_{q}^{\dagger}U]/2$ at leading $p^{2}$ order in the chiral expansion, and terms such as $\mathcal{O}_{\rm KM}$ at higher order. The second is a sum of operators with at most two derivatives generated at the scale $m_{\varphi}$.
  • One operator first appears at $p^{4}$ order in the chiral expansion
@@
In this language, the strong CP problem is solved because the operators are invariant under a global PQ rephasing of $U$, that arises when $\tilde{M}_{q}$ has at least a zero eigenvalue.
The Murayama model is now encoded in the coefficients $C_{i}$. Integrating out the PQ scalar at tree level gives
@@
At loop level one expects order unity corrections. In usual chiral perturbation theory, the coefficient $C_{1}$ would be chirally suppressed by $m_{\pi}^{2}/(4\pi f_{\pi})^{2}$, giving a suppressed contribution to pion masses. However the Murayama theory predicts a larger $C_{1}$ because ${\cal O}_{1}^{\rm PQ}$ substitutes quark masses. So the theory breaks the usual chiral expansion in powers of $p^{2}$. Unlike the LO $p^{2}$ operator, $p^{4}$ operators (including ${\cal O}_{1}^{\rm PQ}$) contribute to pion masses in a way that breaks SU(2) isospin.
@@ 4.1 Pion masses
$m_{\pi_{\pm}}^{2}$ $\approx$ $B_{0}\tilde{m}\bigg(1+\frac{x_{\rm KM}m_{s}}{2}\bigg)+\frac{2C_{1}(\kappa_{u}n_{u}+\kappa_{d}n_{d})^{2}}{f_{\pi}^{2}}.$ (45)
where $\tilde{m}\equiv m_{u}(1-n_{u})+m_{d}(1-n_{d})$. We see that $C_{1}$ breaks isospin at the level of the mass spectrum. In agreement with our previous results, $m_{\pi^{\pm}}^{2}=B_{0}(m_{u}+m_{d})$ (plus KM corrections) keeps its standard QCD value for $C_{1}$ as in eq. (43). The pion mass difference is
@@
The Kaplan-Manohar ambiguity can help only partially, as (despite some similarity) ${\cal O}_{1}^{\rm PQ}$ differs from the Kaplan-Manohar operator in eq. (20).
Isospin SU(2) can perhaps be recovered by adding a SU(2) multiplet of extra scalars at the QCD scale, but the electrically charged components would be excluded.
Loop corrections can become more significant, since $C_{1}$ violates the chiral expansion. Nevertheless, loop corrections cannot cancel the tree-level effect, as long as one operator is dominant.
One operator dominates, so its coefficient cannot be tuned to zero in an extended theory with more GeV-scale PQ scalars. One could consider an extension of the theory where extra operators are added just to tune the pion masses to their observed values. The EFT formalism shows however that this leaves other problems, in pion interactions.
@@ 4.2 Pion interactions
which vanishes in the soft pion limits, a feature known as Adler’s zero. The $s,t,u$ terms arise from the pion kinetic term, while the constant terms arise from the pion mass term. This term changes if the higher-order chiral operator contributes to pion masses. Indeed, in the same limit, ${\cal O}_{1}^{\rm PQ}$ gives a different isosin-breaking function of pions:
@@ 5 Light invisible axion
A heavy axion resembling a QCD meson would be a new solution to the strong CP problem, but we could not solve its problems. We conclude by considering the same theory in the limiting regime where the axion is much lighter than the QCD scale, so that the EFT of section 4 is no longer applicable. As $f_{\varphi}$ becomes large, the pseudo-scalar component $a$ (and its scalar companion $\sigma$) become lighter, producing an invisible axion scenario with the special property that QCD itself breaks the PQ symmetry, leading to a non-standard cosmological history. In this region of the parameter space the masses of the two scalars read (in the following $n_{q}\leq 1$)
@@
where the former can be obtained from diagonalizing the pseudoscalar mass matrix in the limit $f_{\varphi}\gg f_{\pi}$. In this regime the canonically normalized axion field can be identified with $a$. One may then use the standard axion mass relation [16], expressed in terms of the axion decay constant $f_{a}$, to determine the matching condition
@@
A light axion is subject to the usual astrophysical constraints $f_{\varphi}\gtrsim 10^{8}$ GeV [6]. Reproducing the desired quark masses needs small Yukawa couplings $\kappa_{q}$,
@@
The assumption that PQ breaking is triggered by the QCD condensate has novel cosmological consequences, but also a novel problem. Getting $f_{\varphi}\gg f_{\pi}$ from the small QCD-induced PQ breaking needs a small scalar quartic. So both scalars $a$ and $\sigma$ become extremely light, in a comparable way. This can be seen applying the inequality of eq. (12) to the mass relations of eq. (53), obtaining
$m^{2}_{\sigma}=\frac{m_{u}+m_{d}}{m_{d}}m^{2}_{a}+2\lambda f^{2}_{\varphi}\lesssim\frac{m_{u}+m_{d}}{m_{d}}m^{2}_{a}+10^{-4}\,\text{GeV}^{2}\left(\frac{\text{GeV}}{f_{\varphi}}\right)^{2},$ (56)
where we again focused on the model where only $m_{u}$ arises from PQ breaking. This eventually implies the upper bound $m_{\sigma}\lesssim 2m_{a}$.
@@
The scalar coupling to nucleons induces a Yukawa-like non-relativistic potential (not spin-suppressed). For sufficiently light scalars (in particular in the sub-eV range), this new force can compete with gravity and is therefore tightly constrained by tests of the inverse-square law and searches for equivalence-principle violation (see e.g. [18]). As a representative benchmark, current limits translate into $g_{\sigma N}\lesssim 10^{-13}$ for $m_{\sigma}\approx 1~\text{eV}$ (near the edge of astrophysical constraints), tightening to $g_{\sigma N}\lesssim 10^{-24}$ for $m_{\sigma}\approx 10^{-16}~\text{eV}$, basically ruling out the whole parameter space in the light invisible axion limit. Similar results are obtained in models with $n_{q}=2$, in which the PQ symmetry breaking is triggered by the QCD quark condensate by turning the sign of the $|\varphi|^{2}$ mass term. Also in this case the radial mode cannot be decoupled from the axion and fifth-force constraints exclude this scenario.
Therefore the only viable option appears to add a larger source of spontaneous PQ breaking (such as a larger $m_{\varphi}^{2}<0$), recovering a usual DFSZ-like scenario.
@@ 6 Conclusions
We critically examined the possibility proposed in [1] that the axion might be hidden within hadronic resonances. The novel idea is to introduce a complex scalar field charged under a PQ symmetry, whose breaking is triggered by the QCD chiral condensate. In this way, the up-quark mass term is dynamically generated after confinement, reviving the logic of the $m_{u}=0$ solution to the strong CP problem. If realized in nature, this mechanism would imply that the new scalar degrees of freedom are already observed among known hadrons, with the $\eta(1295)$ being a plausible candidate. However the model predicts an order unity splitting among charged and neutral pions [1]. We confirm this problem, fixing order unity factors. Avoiding this problem relying on the Kaplan-Manohar ambiguity to set $m_{u}\ll m_{d}$ is almost as excluded as using it to more simply set $m_{u}=0$.
Motivated by the hope of avoiding this phenomenological exclusion, we considered uncertainties in pion masses and explored several extensions of the minimal setup. Nevertheless, despite some partial success, we find that this entire class of models is strongly constrained, and in practice excluded, by low-energy mesonic observables. The underlying difficulty can be formulated as a general obstruction. If the new scalar degree(s) of freedom are heavy enough to be hidden among hadronic resonances, they can be integrated out, leading to a modified chiral effective theory for the mesons. The presence of the PQ symmetry introduces a new spurion that explicitly breaks isospin. As a consequence, one novel dominant operator appears in the chiral Lagrangian ($\mathcal{O}_{1}^{\rm PQ}$ in eq. (41)), with a coefficient that must make its effects comparable to those of usual chiral operator arising from quark masses. This alternative PQ-symmetric operator is problematic. At quadratic order in pions, it breaks the mass degeneracy among them. At quartic order this operator induces order-one deviations from the well-tested low-energy pion-pion scattering amplitudes.

### v2

@@ 3.4 Model with axion coupled to the strange quark
$m_{\varphi}=\left(\frac{B_{0}f_{\pi}^{2}}{2m_{s}}\right)^{1/2}\kappa_{s}\approx 0.36~\text{GeV}\,\kappa_{s}$ (31)
@@
Extra tentative models with multiple PQ scalars are discussed in appendix A. One might even consider the possibility of adding a PQ-neutral scalar just to make the $\pi^{0}$ heavier. In section 4 we discuss the general reason that prevents realizing a viable model.
@@ 3.5 Loop corrections to pion masses
All results receive corrections that become large if $\kappa$ is large. Here we attempt an estimate. One-loop corrections to pion masses directly affect the dominant diagonal entries of the meson mass matrix in eq. (16). So there could be a controllable regime where the one-loop corrections cancel the too large tree-level correction to $m_{\pi^{0}}-m_{\pi^{\pm}}$, arising from off-diagonal entries in the meson mass matrix. In appendix B we discuss how the pion splitting could be precisely computed using QCD data in the regime where $\kappa$ is perturbative. We here heuristically derive a similar result by computing the renormalization of the meson masses in the chiral Lagrangian effective theory, including $\varphi$. In addition to isospin-symmetric vertices, it contains a $\sigma\pi^{0}\eta^{(\prime)}$ vertex that only affects the $\pi^{0}$ mass. The linear coupling of $\varphi$ to QCD mesons can be written in terms of $\varphi$-dependent meson mass matrix, obtaining the 1-loop contribution as
@@
where latin indices label QCD mesons. The integrand is positive and the integral is logarithmically UV divergent. More importantly, we cannot apply the chiral Lagrangian at large $Q\gtrsim$ GeV so we estimate the contribution with a cut-off at $\Lambda\sim{\rm GeV}$. Assuming $m_{\varphi}\gtrsim\Lambda$ and neglecting for simplicity off-diagonal terms in the mass matrix, we find
@@
This splitting, when extrapolated to $m_{\varphi}>\Lambda$ agrees, including the sign, with the result of the tree level matching eq. (23) upon identifying $\Lambda=4\pi f_{\pi}$. This however should not be considered as a new contribution to the splitting but rather as an estimate of the uncertainty. In the appendix we present a more precise computation that could in principle allow to determine the correction in terms of QCD correlators.
@@ 4 Heavy axion effective chiral theory
In this section we focus on the regime where the new PQ scalars ($a$ and $\sigma$ in the minimal model) are heavy enough that they could be hidden among QCD resonances with GeV mass. This regime needs $n_{s}=0$ and $n_{u,d}\leq 1$ in the language of the previous section. We here exploit the (mild) separation of scales
@@
to integrate out the PQ scalars, obtaining a modified chiral Lagrangian for the light mesons. Namely we capture the effect of the heavy field by matching the full theory of eq. (3) at $m_{\varphi}$ to a new chiral Lagrangian augmented by operators — invariant under PQ — generated by the new scalar dynamics. This allows to exploit the symmetries to constrain the form of extra effects, including loop effects. The low energy degrees of freedom are again the $\Pi$ fields of eq. (8). The $\eta^{\prime}$ is again kept in the effective field theory (EFT) because of its special role for the PQ symmetry. The meson interactions differ from the usual chiral Lagrangian, and even include different isospin-breaking terms. Indeed, at the matching scale, formal invariants under the full symmetry now involve two spurions:
@@
The first term, $\mathscr{L}_{\rm QCD}(\tilde{M}_{q})$, is the chiral Lagrangian with quark masses given by $\tilde{M}_{q}$. As in eq. (6), it includes the term ${f_{\pi}^{2}B_{0}}\operatorname{tr}[\tilde{M}_{q}U^{\dagger}+\tilde{M}_{q}^{\dagger}U]/2$ at leading $p^{2}$ order in the chiral expansion, and terms such as $\mathcal{O}_{\rm KM}$ at higher order. The second is a sum of operators with at most two derivatives generated at the scale $m_{\varphi}$. Focusing on $n_{q}=0,1$,
  • One operator appears at $p^{4}$ order in the chiral expansion
@@
In this language, the strong CP problem is solved because the operators are invariant under a global PQ transformation of $U$, that arises when $\tilde{M}_{q}$ has at least a zero eigenvalue. Therefore the total effective action is PQ invariant up to the anomaly term.
The model [1] is now encoded in the coefficients $C_{i}$. Integrating out the PQ scalar at tree level gives
@@
At loop level one expects sizable corrections in light of the large coupling $\kappa$. In the standard power counting of chiral perturbation theory, the coefficient $C_{1}$ would be suppressed by $m_{\pi}^{2}/(4\pi f_{\pi})^{2}$, giving a small contribution to pion masses. However the Murayama theory predicts a larger $C_{1}$ because ${\cal O}_{1}^{\rm PQ}$ reproduces the effect of up quark mass. So the theory breaks the usual chiral expansion in powers of $p^{2}$. Unlike the leading order $p^{2}$ operator, $p^{4}$ operators (including ${\cal O}_{1}^{\rm PQ}$) contribute to pion masses in a way that breaks SU(2) isospin.
@@ 4.1 Pion masses
$m_{\pi_{\pm}}^{2}$ $\approx$ $B_{0}\tilde{m}\bigg(1+\frac{x_{\rm KM}m_{s}}{2}\bigg)+\frac{2C_{1}(\kappa_{u}n_{u}+\kappa_{d}n_{d})^{2}}{f_{\pi}^{2}},$ (44)
where $\tilde{m}\equiv m_{u}(1-n_{u})+m_{d}(1-n_{d})$. We see that $C_{1}$ breaks isospin at the level of the mass spectrum. For the charged pion one finds the standard result $m_{\pi^{\pm}}^{2}=B_{0}(m_{u}+m_{d})$ (plus KM corrections) for $C_{1}$ as in eq. (42). The pion mass difference is
@@
A few comments are in order. The KM ambiguity can help only partially, as (despite some similarity) ${\cal O}_{1}^{\rm PQ}$ differs from the KM operator in eq. (19). Isospin SU(2) could perhaps be recovered by adding a SU(2) multiplet of extra scalars at the QCD scale, but the electrically charged components would be excluded. Loop corrections can become more significant, since $C_{1}$ violates the chiral expansion. Nevertheless, loop corrections cannot cancel the tree-level effect, as long as one operator is dominant.
As long as one operator dominates, its coefficient cannot be tuned to zero in an extended theory with more GeV-scale PQ scalars. One could consider an extension of the theory where extra operators are added just to tune the pion masses to their observed values. The EFT formalism shows however that this leaves other problems, in pion interactions.
@@ 4.2 Pion interactions
that vanishes in the soft pion limits, a feature known as Adler’s zero. The $s,t,u$ terms arise from the pion kinetic term, while the constant terms arise from the pion mass term. This term changes if the higher-order chiral operator contributes to pion masses. Indeed, in the same limit, ${\cal O}_{1}^{\rm PQ}$ gives a different isospin-breaking function of pions:
@@ 5 Light invisible axion
A heavy axion resembling a QCD meson would be a qualitatively new solution to the strong CP problem, but we could not cure its phenomenological problems. We conclude by considering the same theory in the limiting regime where the axion is much lighter than the QCD scale, so that the EFT of section 4 is no longer applicable. As $f_{\varphi}$ becomes large, the pseudo-scalar component $a$ (and its scalar companion $\sigma$) become lighter, producing an invisible axion scenario with the special property that QCD itself breaks the PQ symmetry, leading to a non-standard cosmological history. In this region of the parameter space the masses of the two scalars read (in the following $n_{q}\leq 1$)
@@
where the former can be obtained from diagonalizing the pseudoscalar mass matrix in the limit $f_{\varphi}\gg f_{\pi}$ and is just the standard invisible QCD axion mass. Indeed in this regime $a$ is just an ordinary QCD axion. One may then use the standard axion mass relation [20], expressed in terms of the axion decay constant $f_{a}$, to determine the matching condition
@@
A light axion is subject to the usual astrophysical constraints. Reproducing the desired quark masses needs small Yukawa couplings $\kappa_{q}$,
@@
The assumption that PQ breaking is triggered by the QCD condensate has novel cosmological consequences, but also a novel problem. Reproducing $f_{\varphi}\gg f_{\pi}$ from the small QCD-induced PQ breaking demands a small scalar quartic. So both scalars $a$ and $\sigma$ become extremely light, in a comparable way. Indeed using eqs. (10)–(13) for $m_{\varphi}^{2}>0$ one can show,
$m^{2}_{a}<\frac{m_{d}}{m_{u}+m_{d}}m_{\sigma}^{2}<3m^{2}_{a},$ (55)
where we again focused on the model where only $m_{u}$ arises from PQ breaking.
@@
The scalar coupling to nucleons induces a Yukawa-like non-relativistic potential (not spin-suppressed). For sufficiently light scalars (in particular in the sub-eV range), this new force can compete with gravity and is therefore tightly constrained by tests of the inverse-square law and searches for equivalence-principle violation (see e.g. [22]). As a representative benchmark, current limits translate into $g_{\sigma N}\lesssim 10^{-13}$ for $m_{\sigma}\approx 1~\text{eV}$ (near the edge of astrophysical constraints [23]), tightening to $g_{\sigma N}\lesssim 10^{-24}$ for $m_{\sigma}\approx 10^{-16}~\text{eV}$, basically ruling out the whole parameter space in the light invisible axion limit. Similar results are obtained in models with $n_{q}=2$, in which the PQ symmetry breaking is triggered by the QCD quark condensate by turning the sign of the $|\varphi|^{2}$ mass term. Also in this case the radial mode cannot be decoupled from the axion and fifth-force constraints exclude this scenario.
Therefore the only viable option appears to add another source of spontaneous PQ breaking (such as $m_{\varphi}^{2}<0$), recovering a DFSZ-like scenario.
@@ 6 Conclusions
We critically examined the possibility proposed in [1] that the axion might be hidden within hadronic resonances. The novel idea is to introduce a complex scalar field charged under a PQ symmetry, whose breaking is triggered by the QCD chiral condensate. In this way, the up-quark mass term is dynamically generated after confinement, reviving the logic of the $m_{u}=0$ solution to the strong CP problem. If realized in nature, this mechanism would imply that the new scalar degrees of freedom are already observed among known hadrons, with the $\eta(1295)$ being a plausible candidate for the pseudo-scalar component. However the model predicts an order unity splitting among charged and neutral pions [1]. We confirm this problem, fixing order unity factors. Avoiding this problem relying on the Kaplan-Manohar ambiguity to set $m_{u}\ll m_{d}$ is almost as excluded as using it to more simply set $m_{u}=0$.
Motivated by the hope of avoiding this phenomenological exclusion, we considered uncertainties in pion masses and explored several extensions of the minimal setup. Nevertheless, despite some partial success, we find that this entire class of models is strongly constrained, and in practice excluded, by low-energy mesonic observables. The underlying difficulty can be formulated as a general obstruction. If the new scalar degree(s) of freedom are heavy enough to be hidden among hadronic resonances, they can be integrated out, leading to a modified chiral effective theory for the mesons. The presence of the PQ symmetry introduces a new spurion that explicitly breaks isospin. As a consequence, one novel dominant operator appears in the chiral Lagrangian ($\mathcal{O}_{1}^{\rm PQ}$ in eq. (40)), with a large coefficient such that its effects are comparable to those of usual chiral operators arising from quark masses. This alternative PQ-symmetric operator is problematic. At quadratic order in pions, it breaks the mass degeneracy among them. At quartic order this operator induces order-one deviations from the well-tested low-energy pion-pion scattering amplitudes.
@@ Appendix A Renormalizable models with more PQ scalars
Appendix B Effective pion potential
We here show how loop corrections to the pion mass splitting can be computed in the perturbative regime. This also provides an estimate of the splitting in the relevant strong coupling regime discussed in the paper.
The computation is analogous to the electromagnetic splitting of pions using form factors. Following [24], we can compute the effective potential of pions and extract the $\pi^{\pm}-\pi^{0}$ splitting at 1-loop. In the limit where the couplings $\kappa_{q}$ are treated as small perturbations, the effective action for $\varphi$ in the background of $\pi$’s can be computed by writing down the most general action consistent with the symmetry of the chiral Lagrangian. To do so we upgrade $\kappa_{q}\varphi$ to a field $\Phi$ transforming $\Phi\to L\Phi R^{\dagger}$ under the chiral symmetry. Treating $\Phi$ as an external source to quadratic order and in the chiral limit, the effective action in the background of pions is
$\mathscr{L}_{\rm eff}=\Pi_{A}(p)\operatorname{tr}[\Phi\Phi^{\dagger}]+\Pi_{B}(p)\operatorname{tr}[\Phi U^{\dagger}]\operatorname{tr}[\Phi^{\dagger}U]+\Pi_{0}\operatorname{tr}[\Phi U^{\dagger}+\text{h.c.}]\,.$ (70)
The form factors $\Pi_{A,B}(p)$ are determined by the two point function of the QCD operators $\mathcal{O}_{LR}^{ij}\equiv\bar{q}_{L}^{i}q_{R}^{j}$. Exploiting the SU($N_{F}$) flavour symmetry and parity we decompose $\mathcal{O}_{LR}^{ij}$ into irreps as,
$\mathcal{O}_{LR}^{ij}=\frac{\delta^{ij}}{\sqrt{2N_{F}}}\mathcal{O}_{\bf 1}+T^{a}_{ij}\mathcal{O}_{\bf Adj}^{a}\,,\qquad\mathcal{O}_{\mathbf{1}}=(\mathcal{O}_{\mathbf{1},S}-i\mathcal{O}_{\mathbf{1},P})\,,\qquad\mathcal{O}_{\bf Adj}^{a}=(\mathcal{O}_{\mathbf{{\bf Adj}},S}^{a}-i\mathcal{O}_{\mathbf{{\bf Adj}},P}^{a})\,$, (71)
where $T^{a}$ are the generator of the adjoint of SU($N_{F}$). For the relevant case $N_{F}=3$ but we write formulas for general number of flavors. The iso-spin channels are the singlet and the octet with complex operators $\mathcal{O}_{(\mathbf{1},\mathbf{{\bf Adj}})}$. In eq. (71) they have been written as the sum of CP-even (S) and CP-odd (P) operators. Explicitly the $S$ operators are $\mathcal{O}_{\mathbf{1},S}\equiv\sqrt{2/N_{F}}\bar{q}q$ and the $P$ are $\mathcal{O}_{\mathbf{1},P}\equiv\sqrt{2/N_{F}}i\bar{q}\gamma_{5}q$, and similarly for the channel in the adjoint. Such operators interpolate with QCD states with the same symmetry properties. We are interested in the momentum space transform of correlators $\langle\mathcal{O}(x)\mathcal{O}(0)\rangle$, given by
$\langle\mathcal{O}_{\mathbf{1},(S,P)}(p)\mathcal{O}_{\mathbf{1},(S,P)}(-p)\rangle=i\Pi_{\mathbf{1},(S,P)}(p)\,,\quad\langle\mathcal{O}_{\mathbf{{\bf Adj}},(S,P)}^{a}(p)\mathcal{O}_{\mathbf{{\bf Adj}},(S,P)}^{b}(-p)\rangle=i\Pi_{\mathbf{{\bf Adj}},(S,P)}(p)\delta^{ab}\,.$ (72)
These correlators only depend on the QCD dynamics and are extracted in the chiral limit. While the general form of the 2-point functions depends on the details of QCD, the situation simplifies in the large $N$ limit where they become a sum of single poles. Moreover the contribution of Nambu-Goldstone bosons is determined by the chiral Lagrangian. We thus have,
$\Pi_{\mathbf{1},P}=\frac{f_{\eta^{\prime}}^{2}B_{0}^{2}}{p^{2}-m_{\eta^{\prime}}^{2}}+\cdots\,,\quad\quad\Pi_{\mathbf{{\bf Adj}},P}=\frac{f_{\pi}^{2}B_{0}^{2}}{p^{2}-m_{\pi}^{2}}+\cdots\,,\quad\Pi_{\mathbf{1},S}=\frac{f_{\sigma_{0}}^{2}B_{0}^{2}}{p^{2}-m_{\sigma_{0}}^{2}}+\cdots\,,\quad\quad\Pi_{\mathbf{{\bf Adj}},S}=\frac{f_{\sigma_{8}}^{2}B_{0}^{2}}{p^{2}-m_{\sigma_{8}}^{2}}+\cdots.$ (73)
In real QCD $f_{\eta^{\prime}}\approx 1.3f_{\pi}$ and the lightest scalar resonances can be identified with $f_{0}(500)$ and isospin triplet $a_{0}(980)$ whose decay constants are not known but are expected to be of order $f_{\pi}$. The form of $\Pi^{\prime}$s is easily understood: the leading contribution to each form factor arises from the exchange of QCD single-particle states with the same quantum numbers of the source, namely, singlet $\eta^{\prime}$, meson octet, single $\sigma_{0}$ resonance and octet of CP-even scalars. The dots represent contributions from heavier resonances and multiparticle states.
The form factors $\Pi_{A,B}$ in eq. (70) are determined by QCD correlation functions above. This can be derived by matching $\Pi_{\bf 1}$ and $\Pi_{\bf Adj}$ to the Lagrangian in eq. (70) on the vacuum $U=1$. Parametrizing $\Phi=\Phi_{1}/\sqrt{2N_{F}}+\Phi_{\bf Adj}^{a}T^{a}$ with $\operatorname{tr}[T^{a}T^{b}]=\delta^{ab}/2$, we obtain,
$\Pi_{A}=-\frac{\Pi_{\mathbf{{\bf Adj}},S}+\Pi_{\mathbf{{\bf Adj}},P}}{2}\,,\qquad\Pi_{B}=\frac{\Pi_{\mathbf{{\bf Adj}},S}+\Pi_{\mathbf{{\bf Adj}},P}-\Pi_{\mathbf{1},S}-\Pi_{\mathbf{1},P}}{2N_{F}}\,.$ (74)
This is reminiscent of the VAV-A structure of the electromagnetic splitting [24]. From the chiral Lagrangian we also extract $\Pi_{0}=f_{\pi}^{2}B_{0}/2$.
Assuming $\Pi_{A,B}$ are known we can proceed to the calculation of the effective potential generated by $\varphi$. For this we make $\varphi$ dynamic adding its action, $|\partial_{\mu}\varphi|^{2}-m_{\varphi}^{2}|\varphi|^{2}$, to eq. (72). For simplicity we can work with only pions in the matrix $U$ so that exponential of the matrix can be explicitly written down. We find,
$\begin{split}\mathscr{L}_{\rm eff}&=\varphi(p)\varphi^{*}(-p)\bigg\{(p^{2}-m_{\varphi}^{2})+\Pi_{A}(p)(\kappa_{u}^{2}n_{u}^{2}+\kappa_{d}^{2}n_{d}^{2})\\ &+\Pi_{B}(p)\left[(\kappa_{u}n_{u}+\kappa_{d}n_{d})^{2}\cos^{2}(\pi)+(\kappa_{u}n_{u}-\kappa_{d}n_{d})^{2}\frac{(\pi^{0})^{2}}{\pi^{2}}\sin^{2}(\pi)\right]\bigg\}\\ &+\Pi_{0}\varphi(p)\left[(\kappa_{u}n_{u}+\kappa_{d}n_{d})\cos(\pi)+i(\kappa_{u}n_{u}-\kappa_{d}n_{d})\frac{\pi^{0}}{\pi}\sin(\pi)\right]+\text{h.c.}\end{split}$ (75)
With these expressions we can easily determine the effective potential performing the path integral over $\varphi$. The term linear in $\varphi$ on the third line generates the tree level contribution obtained integrating out $\varphi$ in section 4. Quantum fluctuations produce a correction,
$\begin{split}\Delta V_{\rm eff}(\pi)=&2\times\frac{-i}{2}\int\frac{d^{4}p}{(2\pi)^{4}}\ln\left[p^{2}-m_{\varphi}^{2}+\left(\sum\kappa_{q}^{2}n_{q}^{2}\right)\Pi_{A}(p)+\right.\\ &\left.+\Pi_{B}(p)\left((\kappa_{u}n_{u}+\kappa_{d}n_{d})^{2}\cos^{2}\pi+(\kappa_{u}n_{u}-\kappa_{d}n_{d})^{2}\frac{(\pi^{0})^{2}}{\pi^{2}}\sin^{2}\pi\right)\right]\,.\end{split}$ (76)
Expanding the logarithm, and Wick rotating to Euclidean momenta $p^{2}\to-Q^{2}$, at leading order in $\kappa_{u,d}$ we can extract the contribution to the pion splitting,
$\Delta(m_{\pi^{0}}^{2}-m_{\pi^{\pm}}^{2})\approx(\kappa_{u}n_{u}-\kappa_{d}n_{d})^{2}\times\frac{2}{f_{\pi}^{2}}\int\frac{d^{4}Q}{(2\pi)^{4}}\frac{\Pi_{B}(Q^{2})}{-Q^{2}-m_{\varphi}^{2}}\,.$ (77)
Knowing the full momentum dependence of the form factors one can precisely determine the mass splitting. Similarly to the electromagnetic splitting we can estimate the integral assuming that it is dominated by the lowest resonances. In the limit where $m_{\varphi}$ is heavy we can require that the integral is finite fixing decay constants $f_{\sigma_{0,8}}$. We find,
$\int\frac{d^{4}Q}{(2\pi)^{4}}\Pi_{B}(Q^{2})=-\frac{B_{0}^{2}\left[f_{\eta^{\prime}}^{2}m_{\eta^{\prime}}^{2}(m_{a_{0}}^{2}-m_{\sigma}^{2})\ln\!\left(\frac{m_{\eta^{\prime}}^{2}}{m_{a_{0}}^{2}}\right)+\left(f_{\pi}^{2}m_{a_{0}}^{2}+f_{\eta^{\prime}}^{2}(m_{\eta^{\prime}}^{2}-m_{a_{0}}^{2})\right)m_{\sigma}^{2}\ln\!\left(\frac{m_{\sigma_{0}}^{2}}{m_{a_{0}}^{2}}\right)\right]}{32\pi^{2}\,(m_{a_{0}}^{2}-m_{\sigma_{0}}^{2})\,\,N_{F}}.$ (78)
Numerically using the observed values of the masses and $N_{F}=3$ we find that the correction to the mass of $\pi_{0}$ is negative and smaller than the tree-level contributions. However, there is no parametric suppression in the result so that the contribution remains uncertain especially in the relevant regime $\kappa_{q}\sim 1$.